OpenStudy (anonymous):

in S.A.= 1/2lp+B is p the perimeter of the base? and in what order do i do them in?

OpenStudy (anonymous):

do your own homework.

Directrix (directrix):

This looks as if it might be the formula for the surface area of a regular pyramid. If so, p is the perimeter of the base and L is the slant height of the pyramid. B is the area of the base of the pyramid. Depending upon which polygon is the base of the pyramid, the formula for B changes. When you ask about order, I don't know what you are asking. Think about posting the specific problem on which you are working.

OpenStudy (anonymous):

i just need to know about the p thank you

Directrix (directrix):

Suppose this were the problem. Find the surface area of a square pyramid with a base length of 5 cm, and a slant height of 10 cm. S.A.= 1/2lp+B SA = (1/2) * (10)*p + B The base is a square of side 5. A square has 4 equal sides so the perimeter of the base is 5+5+5+5 = 20. And, B is the area of a square of side 5. Area of square = 5*5 = 25 So, SA = (1/2) * (10)*p + 25 becomes SA = (1/2) * 10 * 20 + 25 SA = 125 square cm I hope this helps.